# Error analysis of an acceleration corrected diffusion approximation of Langevin dynamics with background flow

### v2

@@ 1.1. Overview and main result
The acceleration corrected diffusion approximation (1.10) has been used extensively in the fields of turbulent transport and cloud physics  [27, 6, 26]. Without the noise term, this approximation was first proposed by Maxey in his seminal work [19], via formal asymptotic calculations. This approximation was also used in [20] to study the effective diffusivity of inertial particles (see Eq. 4.10 in  [20]). Their derivation is through formal asymptotic calculations of the backward Kolmogorov equation of (1.4).
@@
Let $p\geq 4$, $W_{t}$ be a standard Brownian motion in $\mathbb{R}^{n}$, $(X^{\varepsilon}_{t},V^{\varepsilon}_{t})$ the solution to the system of SDE (1.4), $Z^{\varepsilon}_{t}$ the solution of (1.10), and that they satisfy (1.12). Suppose (H1)-(H3) are true for $p$. Then, for each $T>0$, there exists $C,\varepsilon_{0}>0$ that depends on $T$ such that for $\varepsilon<\varepsilon_{0}$,
(1.14) $\sup_{t\in[0,T]}\boldsymbol{E}\left|X^{\varepsilon}_{t}-Z^{\varepsilon}_{t}\right|^{p}\leq C(1+T)\varepsilon^{p}.$
Furthermore, for every $\varphi\in C^{\infty}(\mathbb{T}^{n})$, there exists a constant $C$ such that
(1.15) $\sup_{t\in[0,T]}\left|\boldsymbol{E}\left(\varphi(X^{\varepsilon}_{t})-\varphi(Z^{\varepsilon}_{t})\right)\right|\leq C\varepsilon^{2}\,.$
@@
$dX_{t}=(b-\varepsilon(b_{t}+D_{x}b\,b))\,dt+\sqrt{2\varepsilon}\,dW_{t}$ $\varepsilon$ $\varepsilon^{2}$
@@ 1.2. General Fast-Slow Systems
Finding the optimal $\delta$ was posed as a challenging open problem [3, 14]. When restricting the setting to that of the Langevin equation, as a corollary of Theorem 1.1, we are able to improve Kifer and Bahktin’s result for (1.4) to achieve the optimal rate:
@@
Let $W_{t}$ be a standard Brownian motion, $(X^{\varepsilon}_{t},V^{\varepsilon}_{t})$ be solution to the system of SDE (1.4) and let $H^{\varepsilon}_{t}$ be the solution of (1.17). Then, for each $T>0$, there exists a constant $C$ such that
(1.19) $\sup_{t\in[0,T]}\boldsymbol{E}\left|X^{\varepsilon}_{t}-H^{\varepsilon}_{t}\right|^{p}\leq C\varepsilon^{p}\,.$
Furthermore, for every $\varphi\in C^{\infty}(\mathbb{T}^{n})$, there exists $C>0$ such that
(1.20) $\sup_{t\in[0,T]}\left|\boldsymbol{E}\left(\varphi(X^{\varepsilon}_{t})-\varphi(H^{\varepsilon}_{t})\right)\right|\leq C\varepsilon\,.$
@@ 1.3. Long time behavior
However, the stationary distribution of (1.5) (if it exists) may not necessarily be spatially uniform. In fact, one of our initial motivations for this study was to explain the recently made experimental observations of non-uniform stationary concentration fields of passive particles under incompressible flow [7]. A sample Monte-Carlo simulation of the stationary spatial distribution for (1.5) is given in Figure 4. Note that the spatial non-uniformity is greater than three-fold between the minimum and maximum of this plot, even when $\varepsilon$ is small. In fact, even in the limit of $\varepsilon\to 0$, the stationary distribution is not expected to become uniform. Equation (1.9) will not be able to approximate this stationary distribution given our discussion above. In the same figure, we plot the stationary distribution of the acceleration corrected approximation (1.11), which is seen to provide a good approximation to the stationary distribution of (1.5). This then demonstrates the potential utility of (1.11) for studying the long-time behavior of (1.4). This is somewhat surprising given that the difference between (1.11) and (1.9) is only a term proportional to $\varepsilon$.
@@ 2.1. Warm-up
(2.3) $\sup_{t\in[0,T]}\left|\boldsymbol{E}\left[\varphi(X^{\varepsilon}_{t})-\varphi(Z^{\varepsilon}_{t})\right]\right|\leq C\varepsilon^{2}\,.$
@@
(2.9) $\tilde{V}^{\varepsilon}_{t}=V_{0}e^{-t/\varepsilon}+\sqrt{\frac{2}{\varepsilon}}P_{t}=V_{0}e^{-t/\varepsilon}+\sqrt{2\varepsilon}(dW_{t}-dP_{t}).$
@@
(2.14) $\boldsymbol{E}(X^{\varepsilon,i}_{t}X^{\varepsilon,j}_{t}-2\varepsilon W_{t}^{i}W_{t}^{j})=\mathcal{O}(\varepsilon^{2}).$
@@
(2.15) $\boldsymbol{E}\left\lvert X_{t}^{\varepsilon}\right\rvert^{4}\leq C\boldsymbol{E}\varepsilon^{4}|V_{0}|^{4}+C\varepsilon^{2}\boldsymbol{E}(|W_{t}|^{4}+|P_{t}|^{4})\leq C\varepsilon^{2}$
@@
(2.16) $\boldsymbol{E}\left(\varphi(X^{\varepsilon}_{t})-\varepsilon\partial_{ij}^{2}\varphi(0)W^{i}_{t}W^{j}_{t}\right)=\mathcal{O}(\varepsilon^{2}).$
@@ 2.3. Weak estimate
Note that because $b\in C^{\infty}(\mathbb{T}^{n}\times[0,\infty))$, it is true that $F\in C^{\infty}(\mathbb{T}^{n}\times[0,\infty))$. By regularity theory for Fokker-Planck equation (see [28, Theorem 3.2.4] for example), we have for each $k\in\mathbb{N}$, there is a constant $C$, such that
@@ 3. Limiting Langevin Calculus
(3.8) $\begin{split}&\boldsymbol{E}\int_{0}^{T}G_{t}\cdot\sqrt{\frac{2}{\varepsilon}}P_{t}\,dt=-\sqrt{2\varepsilon}\,\boldsymbol{E}\int_{0}^{T}G_{t}\cdot dJ_{t}\\ =&-\sqrt{2\varepsilon}\,\boldsymbol{E}\left(G_{T}\cdot J_{T}\right)+\sqrt{2\varepsilon}\,\boldsymbol{E}\int_{0}^{T}J_{t}\cdot G_{t}^{\prime}\,dt\,.\end{split}$
@@
(3.10) $\begin{split}&\left(\boldsymbol{E}|G_{T}^{2}|\right)^{1/2}\leq\left(\boldsymbol{E}|G_{0}|^{2}\right)^{1/2}+\left(\boldsymbol{E}\left\lvert\int_{0}^{T}G^{\prime}_{t}dt\right\rvert^{2}\right)^{1/2}\\ \leq&\left(\boldsymbol{E}G_{0}^{2}\right)^{1/2}+\sqrt{T}\left(\boldsymbol{E}\int_{0}^{T}\left\lvert G^{\prime}_{t}\right\rvert^{2}\,dt\right)^{1/2}=C_{1}+\sqrt{T}C_{2}.\end{split}$
@@
(3.11) $\begin{split}&\left|\sqrt{2\varepsilon}\,\boldsymbol{E}\int_{0}^{T}J_{t}\cdot G_{t}^{\prime}\,dt\right|\leq\sqrt{2\varepsilon}\int_{0}^{T}\left(\boldsymbol{E}\left\lvert J_{t}\right\rvert^{2}\right)^{1/2}\left(\boldsymbol{E}\left\lvert G_{t}^{\prime}\right\rvert^{2}\right)^{1/2}dt\\ \leq&\varepsilon\sqrt{nT}\left(\boldsymbol{E}\int_{0}^{T}\left\lvert G^{\prime}_{t}\right\rvert^{2}\,dt\right)^{1/2}\leq\varepsilon\sqrt{nT}C_{2}.\end{split}$
@@
(3.16) $\boldsymbol{E}\left|\int_{0}^{t}\tilde{V}^{\varepsilon}_{s}\,ds\right|^{p}\leq C(t\varepsilon^{p/2}+\varepsilon^{p})\,$,
@@
$\leq 4^{p-1}\left(\varepsilon^{p}\boldsymbol{E}\left|\tilde{V}^{\varepsilon}_{0}\right|^{p}\left(1-e^{-t/\varepsilon}\right)^{p}+(2\varepsilon)^{p/2}\boldsymbol{E}|P_{t}|^{p}+(2\varepsilon)^{p/2}\boldsymbol{E}|W_{t}|^{2}\right)$
$\leq C\left(\varepsilon^{p}\boldsymbol{E}\left|\tilde{V}^{\varepsilon}_{0}\right|^{p}\left(1-e^{-t/\varepsilon}\right)^{p}+\varepsilon^{p}\right)+C\varepsilon^{p/2}t\,$,
@@ 6.1. Auxiliary estimates
(6.3) $\left|\boldsymbol{E}\left(P_{T}\cdot\int_{0}^{T}f(t)\left(\Phi({X^{\varepsilon}_{t}},t)-\Phi(\boldsymbol{E}{X^{\varepsilon}_{t}},t)\right)\,dt\right)\right|\leq C\left(1+\sqrt{T}\right)\varepsilon^{5/2}\,.$
@@
(6.6) $\left\lvert\boldsymbol{E}\int_{0}^{T}\tilde{V}^{\varepsilon}_{t}\cdot\int_{0}^{t}f(s)\Phi({X^{\varepsilon}_{s}},s)\,ds\,dt\right\rvert\leq C\varepsilon^{2}\left(1+T^{3/2}\right)\,.$
@@
$\leq\int_{0}^{T}\left(C_{1}\varepsilon(1+t)^{2}e^{(C_{2}-1/\varepsilon)t}+C_{3}(1+\sqrt{t})\varepsilon^{2}\right)\,dt$
$\leq C\varepsilon^{2}(1+T^{3/2})\,.$
@@
(6.9) $\left\lvert\boldsymbol{E}(IB)\right\rvert\leq C\left(1+T^{3/2}\right)\varepsilon^{2}\,$,
@@
$\left\lvert\boldsymbol{E}(IB)\right\rvert\leq C\left(1+T^{3/2}\right)\varepsilon^{2}\,.$
@@
(6.12) $\left\lvert\boldsymbol{E}\int_{0}^{T}\tilde{V}^{\varepsilon}_{t}\cdot\int_{0}^{t}\Phi(X^{\varepsilon}_{s},s)\,ds\,dt\right\rvert\leq C\varepsilon^{2}(1+T^{3/2})$
@@
(6.13) $\leq C\int_{0}^{T}e^{-s/\varepsilon}\varepsilon^{3/2}\,ds\leq C\varepsilon^{3}\,.$
@@
$\left\lvert\boldsymbol{E}(IB)\right\rvert\leq C\left(1+T^{3/2}\right)\varepsilon^{2}\,.$
@@
$\left|\boldsymbol{E}(IC4)-\varepsilon\int_{0}^{T}\Delta_{x}u(X^{\varepsilon}_{t},t)\,dt\right|=C\left(1+\sqrt{T}\right)\varepsilon^{2}\,.$
@@
(6.24) $\left|\boldsymbol{E}(I)\right|\leq C\left(1+T^{3/2}\right)\varepsilon^{2}\,$,
@@
$\leq\left\lvert\boldsymbol{E}(I)\right\rvert+\left\lvert\boldsymbol{E}(II)\right\rvert\leq C(1+T^{3/2})\varepsilon^{2}\,.$
@@
Note that $\left\lVert\mathcal{T}b\right\rVert_{L^{\infty}(\mathbb{T}^{n}\times[0,T];\mathbb{R}^{n})}\leq C_{T}<\infty$. Estimate (1.19) is a consequence of this observation and Gronwall inequality.

### v3

@@ 1.1. Overview and main result
The acceleration corrected diffusion approximation (1.10) has been used extensively in the fields of turbulent transport and cloud physics  [30, 7, 29]. Recently, a few works started to notice that the appearance of the accelerated correction, though small, may lead to new challenges in homogenization and enhanced dissipation problems that come from inertial particles [6, 27] . However, to the best of our knowledge, rigorous quantitative treatment of the effect of this correction term remains very limited, even non-existent.
Without the noise term, this approximation was first proposed by Maxey in his seminal work [21], via formal asymptotic calculations. This approximation was also used in [22] to study the effective diffusivity of inertial particles (see Eq. 4.10 in  [22]). Their derivation is through formal asymptotic calculations of the backward Kolmogorov equation of (1.4).
@@
Fix $T>0$ and $p\geq 4$. Let $W_{t}$ be a standard Brownian motion in $\mathbb{R}^{n}$, $(X^{\varepsilon}_{t},V^{\varepsilon}_{t})$ the solution to the system of SDE (1.4), $Z^{\varepsilon}_{t}$ the solution of (1.10), and that they satisfy (1.12). Suppose (H1)-(H3) are true for $p$. Then, there exist constants $C_{T}$ and $\varepsilon_{0}>0$ depending on $T$ such that for $\varepsilon<\varepsilon_{0}$,
(1.14) $\sup_{t\in[0,T]}\boldsymbol{E}\left|X^{\varepsilon}_{t}-Z^{\varepsilon}_{t}\right|^{p}\leq C_{T}\varepsilon^{p}.$
Furthermore, for every $\varphi\in C^{\infty}(\mathbb{T}^{n})$, there exists $C_{T}>0$,
(1.15) $\sup_{t\in[0,T]}\left|\boldsymbol{E}\left(\varphi(X^{\varepsilon}_{t})-\varphi(Z^{\varepsilon}_{t})\right)\right|\leq C_{T}\varepsilon^{2}\,.$
@@
$dX_{t}=(b-\varepsilon(\partial_{t}b+D_{x}b\,b))\,dt+\sqrt{2\varepsilon}\,dW_{t}$ $\varepsilon$ $\varepsilon^{2}$
@@ 1.2. General Fast-Slow Systems
Finding the optimal $\delta$ was posed as a challenging open problem [3, 15, 16]. When restricting the setting to that of the Langevin equation, as a corollary of Theorem 1.1, we are able to improve Kifer and Bahktin’s result for (1.4) to achieve the optimal rate:
@@
Fix $T>0$ and $p\geq 4$. Let $W_{t}$ be a standard Brownian motion, $(X^{\varepsilon}_{t},V^{\varepsilon}_{t})$ be solution to the system of SDE (1.4) and let $H^{\varepsilon}_{t}$ be the solution of (1.17). Then, there exist constants $C_{T}$ and $\varepsilon_{0}>0$ depending on $T$ such that for every $\varepsilon<\varepsilon_{0}$,
(1.19) $\sup_{t\in[0,T]}\boldsymbol{E}\left|X^{\varepsilon}_{t}-H^{\varepsilon}_{t}\right|^{p}\leq C_{T}\varepsilon^{p}\,.$
Furthermore, for every $\varphi\in C^{\infty}(\mathbb{T}^{n})$, there exists $C_{T}>0$ such that
(1.20) $\sup_{t\in[0,T]}\left|\boldsymbol{E}\left(\varphi(X^{\varepsilon}_{t})-\varphi(H^{\varepsilon}_{t})\right)\right|\leq C_{T}\varepsilon\,.$
Remark 1.3.
We stress that the estimate (1.20) is of order $\mathcal{O}(\varepsilon)$ (not $\mathcal{O}(\varepsilon^{2})$ as in (1.15)) because it lacks the correction term $\varepsilon(\partial_{t}b+D_{x}bb)$ of the drift $b$. We refer the reader to Remark (5.1) for further details. As discussed in Section 1.3 below, the difference between (1.10) and (1.17) is not only quantitative but also, more strikingly, qualitative.
@@ 1.3. Long time behavior
However, the stationary distribution of (1.5) (if it exists) may not necessarily be spatially uniform. In fact, one of our initial motivations for this study was to explain the recently made experimental observations of non-uniform stationary concentration fields of passive particles under incompressible flow [8].
A sample Monte-Carlo simulation of the stationary spatial distribution for (1.5) is given in Figure 1. Note that the spatial non-uniformity is greater than three-fold between the minimum and maximum of this plot, even when $\varepsilon$ is small. In fact, even in the limit of $\varepsilon\to 0$, the stationary distribution is not expected to become uniform. Equation (1.9) will not be able to approximate this stationary distribution given our discussion above. In the same figure, we plot the stationary distribution of the acceleration corrected approximation (1.11), which is seen to provide a good approximation to the stationary distribution of (1.5). This then demonstrates the potential utility of (1.11) for studying the long-time behavior of (1.4). This is somewhat surprising given that the difference between (1.11) and (1.9) is only a term proportional to $\varepsilon$.
@@ 2.1. Warm-up
(2.3) $\sup_{t\in[0,T]}\left|\boldsymbol{E}\left[\varphi(X^{\varepsilon}_{t})-\varphi(Z^{\varepsilon}_{t})\right]\right|\leq C\varepsilon^{2}T^{2}\,.$
@@
(2.9) $\tilde{V}^{\varepsilon}_{t}\,dt=V_{0}e^{-t/\varepsilon}\,dt+\sqrt{\frac{2}{\varepsilon}}P_{t}\,dt=V_{0}e^{-t/\varepsilon}\,dt+\sqrt{2\varepsilon}(dW_{t}-dP_{t}).$
@@
(2.14) $\left|\boldsymbol{E}(X^{\varepsilon,i}_{t}X^{\varepsilon,j}_{t}-2\varepsilon W_{t}^{i}W_{t}^{j})\right|=C\varepsilon^{2}.$
@@
(2.15) $\boldsymbol{E}\left\lvert X_{t}^{\varepsilon}\right\rvert^{4}\leq C\boldsymbol{E}\varepsilon^{4}|V_{0}|^{4}+C\varepsilon^{2}\boldsymbol{E}(|W_{t}|^{4}+|P_{t}|^{4})\leq C\varepsilon^{2}t^{2}$
@@
(2.16) $\left|\boldsymbol{E}\left(\varphi(X^{\varepsilon}_{t})-\varepsilon\partial_{ij}^{2}\varphi(0)W^{i}_{t}W^{j}_{t}\right)\right|\leq C\varepsilon^{2}t^{2}.$
@@ 2.3. Weak estimate
Note that because $b\in C^{\infty}(\mathbb{T}^{n}\times[0,\infty))$, it is true that $F\in C^{\infty}(\mathbb{T}^{n}\times[0,\infty))$. By regularity theory for Fokker-Planck equation (see [31, Theorem 3.2.4] for example), we have for each $k\in\mathbb{N}$, there exist a constant $C=C_{T}>0$, such that
@@ 3. Limiting Langevin Calculus
(3.8) $\begin{split}&\boldsymbol{E}\int_{0}^{T}G_{t}\cdot\sqrt{\frac{2}{\varepsilon}}J_{t}\,dt=-\sqrt{2\varepsilon}\,\boldsymbol{E}\int_{0}^{T}G_{t}\cdot dJ_{t}\\ &=-\sqrt{2\varepsilon}\,\boldsymbol{E}\left(G_{T}\cdot J_{T}\right)+\sqrt{2\varepsilon}\,\boldsymbol{E}\int_{0}^{T}J_{t}\cdot G_{t}^{\prime}\,dt\,.\end{split}$
@@
(3.10) $\begin{split}&\left(\boldsymbol{E}|G_{T}^{2}|\right)^{1/2}\leq\left(\boldsymbol{E}|G_{0}|^{2}\right)^{1/2}+\left(\boldsymbol{E}\left\lvert\int_{0}^{T}G^{\prime}_{t}dt\right\rvert^{2}\right)^{1/2}\\ &\leq\left(\boldsymbol{E}G_{0}^{2}\right)^{1/2}+\sqrt{T}\left(\boldsymbol{E}\int_{0}^{T}\left\lvert G^{\prime}_{t}\right\rvert^{2}\,dt\right)^{1/2}=C_{1}+\sqrt{T}C_{2}.\end{split}$
@@
(3.11) $\begin{split}&\left|\sqrt{2\varepsilon}\,\boldsymbol{E}\int_{0}^{T}J_{t}\cdot G_{t}^{\prime}\,dt\right|\leq\sqrt{2\varepsilon}\int_{0}^{T}\left(\boldsymbol{E}\left\lvert J_{t}\right\rvert^{2}\right)^{1/2}\left(\boldsymbol{E}\left\lvert G_{t}^{\prime}\right\rvert^{2}\right)^{1/2}dt\\ &\leq\varepsilon\sqrt{nT}\left(\boldsymbol{E}\int_{0}^{T}\left\lvert G^{\prime}_{t}\right\rvert^{2}\,dt\right)^{1/2}\leq\varepsilon\sqrt{nT}C_{2}.\end{split}$
@@
(3.16) $\boldsymbol{E}\left|\int_{0}^{t}\tilde{V}^{\varepsilon}_{s}\,ds\right|^{p}\leq C(t^{p/2}\varepsilon^{p/2}+\varepsilon^{p})\,$,
@@
$\leq 4^{p-1}\left(\varepsilon^{p}\boldsymbol{E}\left|\tilde{V}^{\varepsilon}_{0}\right|^{p}\left(1-e^{-t/\varepsilon}\right)^{p}+(2\varepsilon)^{p/2}\boldsymbol{E}|P_{t}|^{p}+(2\varepsilon)^{p/2}\boldsymbol{E}|W_{t}|^{p}\right)$
$\leq C\left(\varepsilon^{p}\boldsymbol{E}\left|\tilde{V}^{\varepsilon}_{0}\right|^{p}\left(1-e^{-t/\varepsilon}\right)^{p}+\varepsilon^{p}\right)+C\varepsilon^{p/2}t^{p/2}\,$,
@@ 6.1. Auxiliary estimates
(6.3) $\left|\int_{0}^{T}\boldsymbol{E}\left(P_{t}\cdot\int_{0}^{t}f(s)\left(\Phi({X^{\varepsilon}_{s}},s)-\Phi(\boldsymbol{E}{X^{\varepsilon}_{s}},s)\right)\,ds\right)\,dt\right|\leq C\left(1+\sqrt{T}\right)\varepsilon^{5/2}\,.$
@@
(6.6) $\left\lvert\boldsymbol{E}\int_{0}^{T}\tilde{V}^{\varepsilon}_{t}\cdot\int_{0}^{t}f(s)\Phi({X^{\varepsilon}_{s}},s)\,ds\,dt\right\rvert\leq C\varepsilon^{2}\left(1+\sqrt{T}\right)\,.$
@@
$\leq\int_{0}^{T}\left(C_{1}\varepsilon(1+t)^{2}e^{(C_{2}-1/\varepsilon)t}\right)\,dt+C_{3}(1+\sqrt{T})\varepsilon^{2}$
$\leq C\varepsilon^{2}(1+\sqrt{T})\,.$
@@
(6.9) $\left\lvert\boldsymbol{E}(IB)\right\rvert\leq C\left(1+T\right)\varepsilon^{2}\,$,
@@
$\left\lvert\boldsymbol{E}(IB)\right\rvert\leq C\left(1+T\right)\varepsilon^{2}\,.$
@@
(6.12) $\left\lvert\boldsymbol{E}\int_{0}^{T}\tilde{V}^{\varepsilon}_{t}\cdot\int_{0}^{t}\Phi(X^{\varepsilon}_{s},s)\,ds\,dt\right\rvert\leq C\varepsilon^{2}(1+\sqrt{T})$
@@
(6.13) $\leq C\int_{0}^{T}e^{-s/\varepsilon}\varepsilon^{2}\,ds\leq C\varepsilon^{3}\,.$
@@
$\left\lvert\boldsymbol{E}(IB)\right\rvert\leq C\left(1+T\right)\varepsilon^{2}\,.$
@@
(6.24) $\left|\boldsymbol{E}(IC4)-\varepsilon\int_{0}^{T}\Delta_{x}u(X^{\varepsilon}_{t},t)\,dt\right|=C\left(1+T\right)\varepsilon^{2}\,.$
@@
(6.25) $\left|\boldsymbol{E}(I)\right|\leq C\left(1+T\right)\varepsilon^{2}\,$,
@@
$\leq\left\lvert\boldsymbol{E}(I)\right\rvert+\left\lvert\boldsymbol{E}(II)\right\rvert\leq C(1+T)\varepsilon^{2}\,.$
@@
Note that $\left\lVert\mathcal{T}b\right\rVert_{L^{\infty}(\mathbb{T}^{n}\times[0,T];\mathbb{R}^{n})}\leq C<\infty$. Estimate (1.19) is a consequence of this observation and Gronwall inequality.